Find the ratio in which the line
3x + y - 9 = 0 divides the line segment joining the points (1,3) and (2,7).

Grade:12th Pass

3 Answers

PRIYANKA TATA ASKIITIAN
13 years ago

to solve this question..first find the line equation frm the two points..on solving u will get it as 4x-y=1-----(1)

now find the intersection point of the two given lines...we get(10/7,33/7)

then u can apply section formula and find the ratio..

Note: if it is asked as a multiple choice question.,u have to find the ratio both internally and externally.

Smile

Mir Noaman Ali IIT-Roorkee
13 years ago

ansr 3:4 internally find the perpendicular distance they will come 3/root 10 and 4/root 10 hence ansr 3:4

Durgesh Prasher
13 years ago

the value of ratio is -L1/L2

        where L1=3[1]+3-9=-3

                 L2=3[2]+7-9=4

                 ratio will 3/4